In elastic collision, when a massive object collides with light body at conditions m1>>m2 and v2= 0m/s, then the change in velocity will be written as
1
Expert's answer
2020-11-20T09:29:43-0500

Conservation of momentum:

"m_1v_1=m_1v_1'+m_2v_2'"

Conservation of energy:


"0.5m_1v_1^2=0.5m_1v_1'^2+0.5m_2v_2'^2"

The change in velocity will be written as


"\Delta v_1=v_1-v_1\frac{m_1-m_2}{m_1+m_2}=\frac{2m_2v_1}{m_1+m_2}"